OpenStudy (anonymous):

has anyone solved ques 2 ..part 2... R1=30Ω, R2=15Ω, R3=15Ω, R4=30Ω,V=15V and dependent source is of 4i..

OpenStudy (anonymous):

wont the above formula give vth =0..??

OpenStudy (anonymous):

please help for the vth part... as per your formula...and i have r1=30 , r2=15,r3=15,r4=30...i will get vth = 0 v....its worong...please help...

OpenStudy (anonymous):

check this out.. Vth=V*((R1/(R1+R3)-(R2/*(R2+R4)) Rth=((R1/(R1+R3))+(R2/(R2+R4)) modifed Vth=VTH(same) modifed Rth=RTh/5

OpenStudy (anonymous):

Sry... Vth=V*((R2/(R2+R4)-(R1/(R3+R1)).. I'm sure of this